OpenStudy (annie96):

Find the slope and y-intercept of the line. y = –x + 7 slope: –7, y-intercept: 1 slope: –1, y-intercept: 7 slope: 1, y-intercept: –7 slope: 7, y-intercept: –1

OpenStudy (anonymous):

y = mx + b m = slope b = y-intercept Your equation is y = -x + 7

OpenStudy (anonymous):

So this is the same as saying y = (-1)x + 7

OpenStudy (annie96):

okay

OpenStudy (anonymous):

So what is the answer?

OpenStudy (annie96):

b

OpenStudy (anonymous):

y = mx + b m = slope b = y-intercept y = (-1)x + 7 m b

OpenStudy (anonymous):